OpenStudy (anonymous):

a person invested rs 500 at a rate of interest 18% for 3 years .what is the compound interest he get?

OpenStudy (anonymous):

rupee rs means

OpenStudy (anonymous):

compound interest = 100 + 18 =118 Right ?

OpenStudy (anonymous):

how ?.no

OpenStudy (anonymous):

500*118*118*118/1000000-500=ans right ah

OpenStudy (anonymous):

If we get the first mony 100% and the compound interest 18% for a year the mony is 100%+18% = 118% OK ?

OpenStudy (anonymous):

no.i cant

OpenStudy (anonymous):

Why ?!

OpenStudy (anonymous):

how 100%

OpenStudy (anonymous):

Because we have all the mony !!!:)

OpenStudy (anonymous):

321-500=answer i think

OpenStudy (anonymous):

compound interest formula?

OpenStudy (anonymous):

c.i=p(1+r/100)^n

OpenStudy (anonymous):

Hey we have compound interes and in this we cant reduse the mony ! We get mony more !

OpenStudy (anonymous):

We dont need to fomula !

OpenStudy (anonymous):

for 3 years

OpenStudy (anonymous):

I answer you : 100%+18=118% Our mony than a year . 100_______________500 118________________x x=118.500/100=590 than a year .

OpenStudy (anonymous):

For one years again : 100+18=118% 100____________________590 118______________________x x=118.590/100=631/5 For 2th year .

OpenStudy (anonymous):

ya.

OpenStudy (anonymous):

For 3th year : 100+18=118% 100______________________631/5 118_______________________x x=118.631/5/100=38229/10

OpenStudy (anonymous):

That was your answer ! Understand ?

OpenStudy (anonymous):

that only i told earlier.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

500*118*118*118/1000000-500=ans right ah

OpenStudy (anonymous):

Yeah !

OpenStudy (anonymous):

Excuse me to it was iong ! I want to you know good !

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."